3. A long wire lying along the x-axis carries a current of 2.00 A in the +x direction. There is a uniform magnetic field present, given by B-0.00+ 0.002j+0.003k, where i, j, k are the unit vectors along the cartesian coordinate axes) in units of Tesla. (a) Calculate the y-component of the magnetic force acting on a segment of wire of length L-20.0 cm. (b) Calculate the z-component of the magnetic force.
Answer #1

\vec {F} = I \vec {L} \times \vec {B}

\vec {F} = 2 (0.2\hat {i})\times (0.001\hat {i}+0.002 \hat {j}+ 0.003\hat{k}) = 0.0008\hat {k} -0.0012\hat {j}

(a)

F0.0012 N

(b)

F_z = 0.0008 \: N